# Ensuring Safety in an Uncertain Environment: Constrained MDPs via Stochastic Thresholds

### v2

@@ H.2 Proof of Theorem 3
Under the same conditions as Theorem 2, with probability at least $1-\delta$, Algorithm 2 under optimistic thresholds guarantees the following regret bounds:
@@
$+\tilde{\mathcal{O}}\left(\sqrt{mSAH^{4}T/\gamma}+\sqrt{m}SAH^{2}\right).$
where $\rho^{\prime}$ is the Slater gap under optimistic thresholds.
@@
$\mathcal{L}(\lambda)=$ $\sum_{t=1}^{T^{\prime}}(f_{t}-f_{\text{opt}})+\sum_{t=1}^{T^{\prime}}y_{t}^{T}\lambda$
$\leq$ $\sum_{t=1}^{T^{\prime}}(y_{t}-\tilde{y}_{t})^{T}\lambda+\sum_{t=1}^{T^{\prime}}\left(\tilde{f}_{t}+\lambda_{t}^{T}\tilde{y}_{t}-f_{\pi^{\star}}-\lambda_{t}^{T}y_{\pi^{\star}}\right)+\sum_{t=1}^{T^{\prime}}(f_{t}-\tilde{f}_{t})+\frac{1}{2\eta_{\lambda}}\sum_{t=1}^{T^{\prime}}\|\tilde{y}_{t}\|^{2}+\frac{\eta_{\lambda}}{2}\|\lambda\|_{2}^{2}$
$\leq$ $\underbrace{\sqrt{\sum_{i=1}^{m}\left(\sum_{t=1}^{T^{\prime}}(y_{t,i}-\tilde{y}_{t,i})\right)^{2}}\|\lambda\|_{2}}_{(a^{\prime})}+\underbrace{\sum_{t=1}^{T^{\prime}}\left(\tilde{f}_{t}+\lambda_{t}^{T}\tilde{y}_{t}-f_{\pi^{\star}}-\lambda_{t}^{T}y_{\pi^{\star}}\right)}_{(b^{\prime})}$
@@
Unlike pessimistic thresholds, for the optimistic thresholds, it holds that $\lambda_{t,i}(\alpha_{i}-\underline{\alpha}_{t,i})\geq 0$ for any constraint $i$ and episode $t$. We then obtain
$\left|\sum_{t=1}^{T^{\prime}}(y_{t,i}-\tilde{y}_{t,i})\right|$ $=\left|\sum_{t=1}^{T^{\prime}}\left(\hat{V}_{\overline{g}_{i}}^{\pi_{t}}(s_{1})-V_{g_{i}}^{\pi_{t}}(s_{1})\right)+\sum_{t=1}^{T^{\prime}}\left(\alpha_{i}-\underline{\alpha}_{t,i}\right)\right|$
$\leq\left|\sum_{t=1}^{T^{\prime}}\left(\hat{V}_{\overline{g}_{i}}^{\pi_{t}}(s_{1})-V_{g_{i}}^{\pi_{t}}(s_{1})\right)\right|+\left|\sum_{t=1}^{T^{\prime}}\left(\alpha_{i}-\underline{\alpha}_{t,i}\right)\right|.$ (18)
@@
$\leq$ $\sqrt{m}\max_{i\in[m]}\left[\left|\sum_{t=1}^{T^{\prime}}\left(\hat{V}_{\overline{g}_{i}}^{\pi_{t}}(s_{1})-V_{g_{i}}^{\pi_{t}}(s_{1})\right)\right|+\left|\sum_{t=1}^{T^{\prime}}(\alpha_{i}-\underline{\alpha}_{t,i})\right|\right]$
$\leq$ $\tilde{\mathcal{O}}\left(\sqrt{mS\mathcal{N}H^{4}T}+(\sqrt{\mathcal{N}}+H+1)\sqrt{m}H^{2}SA+\sqrt{mSAH^{4}T/\gamma}\right).$
@@
$(a^{\prime})\leq\tilde{\mathcal{O}}\left(\sqrt{mS\mathcal{N}H^{4}T}+(\sqrt{\mathcal{N}}+H+1)\sqrt{m}H^{2}SA+\sqrt{mSAH^{4}T/\gamma}\right)\|\lambda\|_{2}.$
@@
$(b^{\prime})=\sum_{t=1}^{T^{\prime}}\left(\tilde{f}_{t}+\lambda_{t}^{T}\tilde{y}_{t}-f_{\pi^{\star}}-\lambda_{t}^{T}y_{\pi^{\star}}\right)\lesssim(1+m\rho^{\prime})H^{2}\sqrt{T}.$
@@
$\sum_{t=1}^{T^{\prime}}(f_{t}-f_{\text{opt}})+\sum_{t=1}^{T^{\prime}}y_{t}^{T}\lambda\lesssim$ $\left(\sqrt{mS\mathcal{N}H^{4}T}+(\sqrt{\mathcal{N}}+H)\sqrt{m}H^{2}SA\right)\|\lambda\|_{2}$
$+\left(\sqrt{mSAH^{4}T/\gamma}+\sqrt{m}SAH^{2}\right)\|\lambda\|_{2}+(1+m\rho^{\prime})H^{2}\sqrt{T}$
@@
$+\rho^{\prime}\left(\sqrt{mSAH^{4}T/\gamma}+\sqrt{m}SAH^{2}\right)+(1+m\rho^{\prime})H^{2}\sqrt{T}$
@@
By Lemmas 23 and 24, we have the following result for any $T^{\prime}\in[T]$: $\max_{i\in[m]}\left[\sum_{t=1}^{T^{\prime}}y_{t}\right]\leq\frac{\theta^{\prime}(T)}{\rho^{\prime}}$. Therefore, we conclude the bound for constraint violation
@@
$+\tilde{\mathcal{O}}\left(\sqrt{mSAH^{4}T/\gamma}+\sqrt{m}SAH^{2}\right).$
@@ Appendix I Constrained Convex Optimisation
    Each $y_{i}(\cdot)$ is concave for $i\in[m]$ and $y(\cdot):=(y_{1}(\cdot),\ldots,y_{m}(\cdot))^{T}$.
@@
    For every $\lambda\in\mathbb{R}_{\geq 0}^{m}$, the minimisation problem $\min_{x\in X}[f(x)+\lambda^{T}y(x)]$ admits an optimal solution.
@@
$q(\lambda)\;=\;\min_{x\in X}\;\mathcal{L}(x,\lambda)\;=\;\min_{x\in X}\Bigl{[}f(x)+\lambda^{T}y(x)\Bigr{]},$
@@
In this Appendix, we present important preparation lemmas from prior work [12, 11].
Lemma 16 ([11]).
@@
$N_{h}^{t-1}(s,a)>\frac{1}{2}\sum_{j<t}q_{h}^{\pi_{j}}(s,a;p)-H\ln\left(\frac{SAH}{\delta^{\prime}}\right),$
then for all $T^{\prime}\leq T$,
$\sum_{t^{\prime}=1}^{T^{\prime}}\sum_{h=1}^{H}\mathbb{E}\left[\sqrt{\frac{1}{N_{h}^{t^{\prime}-1}(s_{h}^{t^{\prime}},a_{h}^{t^{\prime}})\vee 1}}\mid\mathcal{F}_{t^{\prime}-1}\right]\leq\tilde{\mathcal{O}}(\sqrt{SAH^{2}T^{\prime}}+SAH).$
Lemma 18 ([11]).
@@
Lemma 19 (OMD term bound [11]).
@@
Lemma 20 (Dual optimism [11]).
@@
$\tilde{f}_{t}-f_{\text{opt}}\leq-\lambda_{t}^{T}\tilde{y}_{t}+\left(\tilde{f}_{t}+\lambda_{t}^{T}\tilde{y}_{t}-f_{\pi^{\star}}-\lambda_{t}^{T}y_{\pi^{\star}}\right).$
Lemma 21 (On policy errors for truncated policy estimation [11]).
@@
    $\big{|}\overline{l}_{h}^{t}(s,a)-l_{h}(s,a)\big{|}\lesssim\sqrt{\frac{1}{N_{h}^{t-1}(s,a)}}$.
@@
    $\big{|}\overline{p}_{h}^{t}(s^{\prime}\mid s,a)-p_{h}(s^{\prime}\mid s,a)\big{|}\lesssim\sqrt{\frac{p_{h}(s^{\prime}\mid s,a)}{\max\{N_{h}^{t-1}(s,a),1\}}}+\frac{1}{\max\{N_{h}^{t-1}(s,a),1\}}$.
@@
    $N_{h}^{t-1}(s,a)\leq\frac{1}{2}\sum_{j<t}q_{h}^{\pi_{j}}(s,a;p)-H\ln\frac{SAH}{\delta^{\prime}}$.
@@
Lemma 22 (Update rule recursion bound primal-dual [11]).
@@
$\sum_{t=1}^{T^{\prime}}(-\tilde{y}_{t}^{T}\lambda_{t})+\sum_{t=1}^{N}\tilde{y}_{t}^{T}\lambda\leq\frac{\eta_{\lambda}}{2}\|\lambda_{1}-\lambda\|_{2}^{2}+\frac{1}{2\eta_{\lambda}}\sum_{t=1}^{T^{\prime}}\|\tilde{y}_{t}\|^{2}.$
Lemma 23 ([11]).

### v3

@@ H.2 Proof of Theorem 3
Under the same conditions of Theorem 2, with probability $1-\delta$, Algorithm 2 under optimistic thresholds holds the following regret bounds:
@@
$+\tilde{\mathcal{O}}\left(\sqrt{mSAH^{2}T/\gamma}+\sqrt{m}SAH\right).$
@@
$\mathcal{L}(\lambda)=$ $\sum_{t=1}^{T^{\prime}}(f_{t}-f_{\text{opt}})+\sum_{t=1}^{T^{\prime}}y_{t}^{\top}\lambda$
$\leq$ $\sum_{t=1}^{T^{\prime}}(y_{t}-\tilde{y}_{t})^{\top}\lambda+\sum_{t=1}^{T^{\prime}}\left(\tilde{f}_{t}+\lambda_{t}^{\top}\tilde{y}_{t}-f_{\pi^{\star}}-\lambda_{t}^{\top}y_{\pi^{\star}}\right)+\sum_{t=1}^{T^{\prime}}(f_{t}-\tilde{f}_{t})+\frac{1}{2\eta_{\lambda}}\sum_{t=1}^{T^{\prime}}\|\tilde{y}_{t}\|^{2}+\frac{\eta_{\lambda}}{2}\|\lambda\|_{2}^{2}$
$\leq$ $\underbrace{\sqrt{\sum_{i=1}^{m}\left(\sum_{t=1}^{T^{\prime}}(y_{t,i}-\tilde{y}_{t,i})\right)^{2}}\|\lambda\|_{2}}_{(a^{\prime})}+\underbrace{\sum_{t=1}^{T^{\prime}}\left(\tilde{f}_{t}+\lambda_{t}^{\top}\tilde{y}_{t}-f_{\pi^{\star}}-\lambda_{t}^{\top}y_{\pi^{\star}}\right)}_{(b^{\prime})}$
@@
Unlike pessimistic thresholds, for the optimistic thresholds, it holds that $\lambda_{t,i}(\alpha_{i}-\underline{\alpha}_{i}^{t})\geq 0$ for any constraint $i$ and episode $t$. We then obtain
$\left|\sum_{t=1}^{T^{\prime}}(y_{t,i}-\tilde{y}_{t,i})\right|$ $=\left|\sum_{t=1}^{T^{\prime}}\left(\hat{V}_{\overline{g}_{i}}^{\pi_{t}}(s_{1})-V_{g_{i}}^{\pi_{t}}(s_{1})\right)+\sum_{t=1}^{T^{\prime}}\left(\alpha_{i}-\underline{\alpha}_{i}^{t}\right)\right|$
$\leq\left|\sum_{t=1}^{T^{\prime}}\left(\hat{V}_{\overline{g}_{i}}^{\pi_{t}}(s_{1})-V_{g_{i}}^{\pi_{t}}(s_{1})\right)\right|+\left|\sum_{t=1}^{T^{\prime}}\left(\alpha_{i}-\underline{\alpha}_{i}^{t}\right)\right|.$ (17)
@@
$\leq$ $\sqrt{m}\max_{i\in[m]}\left[\left|\sum_{t=1}^{T^{\prime}}\left(\hat{V}_{\overline{g}_{i}}^{\pi_{t}}(s_{1})-V_{g_{i}}^{\pi_{t}}(s_{1})\right)\right|+\left|\sum_{t=1}^{T^{\prime}}(\alpha_{i}-\underline{\alpha}_{i}^{t})\right|\right]$
$\leq$ $\tilde{\mathcal{O}}\left(\sqrt{mS\mathcal{N}H^{4}T}+(\sqrt{\mathcal{N}}+H+1)\sqrt{m}H^{2}SA+\sqrt{mSAH^{2}T/\gamma}\right).$
@@
$(a^{\prime})\leq\tilde{\mathcal{O}}\left(\sqrt{mS\mathcal{N}H^{4}T}+(\sqrt{\mathcal{N}}+H+1)\sqrt{m}H^{2}SA+\sqrt{mSAH^{2}T/\gamma}\right)\|\lambda\|_{2}.$
@@
$(b^{\prime})=\sum_{t=1}^{T^{\prime}}\left(\tilde{f}_{t}+\lambda_{t}^{\top}\tilde{y}_{t}-f_{\pi^{\star}}-\lambda_{t}^{\top}y_{\pi^{\star}}\right)\lesssim(1+m\rho^{\prime})H^{2}\sqrt{T}.$
@@
$\sum_{t=1}^{T^{\prime}}(f_{t}-f_{\text{opt}})+\sum_{t=1}^{T^{\prime}}y_{t}^{\top}\lambda\lesssim$ $\left(\sqrt{mS\mathcal{N}H^{4}T}+(\sqrt{\mathcal{N}}+H)\sqrt{m}H^{2}SA\right)\|\lambda\|_{2}$
$+\left(\sqrt{mSAH^{2}T/\gamma}+\sqrt{m}SAH\right)\|\lambda\|_{2}+(1+m\rho^{\prime})H^{2}\sqrt{T}$
@@
$+\rho^{\prime}\left(\sqrt{mSAH^{2}T/\gamma}+\sqrt{m}SAH\right)+(1+m\rho^{\prime})H^{2}\sqrt{T}$
@@
∎ By Lemmas 22 and 23, we have the following result for any $T^{\prime}\in[T]$: $\max_{i\in[m]}\left[\sum_{t=1}^{T^{\prime}}y_{t}\right]\leq\frac{\theta^{\prime}(T)}{\rho^{\prime}}$. Therefore, we conclude the bound for constraint violation
@@
$+\tilde{\mathcal{O}}\left(\sqrt{mSAH^{2}T/\gamma}+\sqrt{m}SAH\right).$
@@ Appendix I Constrained Convex Optimisation
    Each $y_{i}(\cdot)$ is concave for $i\in[m]$ and $y(\cdot):=(y_{1}(\cdot),\ldots,y_{m}(\cdot))^{\top}$.
@@
    For every $\lambda\in\mathbb{R}_{\geq 0}^{m}$, the minimisation problem $\min_{x\in X}[f(x)+\lambda^{\top}y(x)]$ admits an optimal solution.
@@
$q(\lambda)\;=\;\min_{x\in X}\;\mathcal{L}(x,\lambda)\;=\;\min_{x\in X}\Bigl[f(x)+\lambda^{\top}y(x)\Bigr],$
@@
In this Appendix, we present important preparation lemmas from prior work [efroni2019tight, efroni2020exploration].
Lemma 15 ([efroni2020exploration]).
@@
$N_{h}^{t}(s,a)>\frac{1}{2}\sum_{j<t}q_{h}^{\pi_{j}}(s,a;p)-H\ln\left(\frac{SAH}{\delta^{\prime}}\right),$
then
$\sum_{t=1}^{T}\sum_{h=1}^{H}\mathbb{E}\left[\sqrt{\frac{1}{N_{h}^{t}(s_{h}^{t},a_{h}^{t})\vee 1}}\mid\mathcal{F}_{t-1}\right]\leq\tilde{\mathcal{O}}(\sqrt{SAH^{2}T}+SAH).$
Lemma 17 ([efroni2020exploration]).
@@
Lemma 18 (OMD term bound [efroni2020exploration]).
@@
Lemma 19 (Dual optimism [efroni2020exploration]).
@@
$\tilde{f}_{t}-f_{\text{opt}}\leq-\lambda_{t}^{\top}\tilde{y}_{t}+\left(\tilde{f}_{t}+\lambda_{t}^{\top}\tilde{y}_{t}-f_{\pi^{\star}}-\lambda_{t}^{\top}y_{\pi^{\star}}\right).$
Lemma 20 (On policy errors for truncated policy estimation [efroni2020exploration]).
@@
    $\big|\hat{l}_{h}^{t}(s,a)-l_{h}(s,a)\big|\lesssim\sqrt{\frac{1}{N_{h}^{t}(s,a)}}$.
@@
    $\big|\overline{p}_{h}^{t}(s^{\prime}\mid s,a)-p_{h}(s^{\prime}\mid s,a)\big|\lesssim\sqrt{\frac{p_{h}(s^{\prime}\mid s,a)}{N_{h}^{t}(s,a)\lor 1}}+\frac{1}{N_{h}^{t}(s,a)\lor 1}$.
@@
    $N_{h}^{t}(s,a)\leq\frac{1}{2}\sum_{j<t}q_{h}^{\pi_{t}}(s,a\mid p)-H\ln\frac{SAH}{\delta^{\prime}}$.
@@
Lemma 21 (Update rule recursion bound primal-dual [efroni2020exploration]).
@@
$\sum_{t=1}^{T^{\prime}}(-\tilde{y}_{t}^{\top}\lambda_{t})+\sum_{t=1}^{N}\tilde{y}_{t}^{\top}\lambda\leq\frac{\eta_{\lambda}}{2}\|\lambda_{1}-\lambda\|_{2}^{2}+\frac{1}{2\eta_{\lambda}}\sum_{t=1}^{T^{\prime}}\|\tilde{y}_{t}\|^{2}.$
Lemma 22 ([efroni2020exploration]).